# An Information Theory of Finite Abstractions and their Fundamental Scalability Limits

### v1

@@ 5.1 A fundamental limit on abstracting dynamical systems
Consider a dynamical system $x^{+}=f(x)$ with transition system representation $S=(\mathcal{X},\underset{S}{\rightarrow})$, and let Assumption 1 hold. Let $\xi\in{\mathcal{B}}_{l}^{S}$ be a trajectory of $S$, with $\xi_{0}\sim p_{\xi_{0}}:\mathcal{X}\to\mathbb{R}_{+}$. Consider a measurable partition $\mathcal{Y}$ of $\mathcal{X}$ and an associated abstraction AA, and let $\Omega_{A}=g_{A}(s_{A}(\xi))$, where $s,g$ are given by (5) and (6). Consider an encoder-decoder pair $(s_{q_{A}},g_{q_{A}})$, where $s_{q_{A}}(\xi)=g_{A}(s_{A}(\xi))$ and $g_{q_{A}}(z)=x_{c}(z)$, where $x_{c}(z):=\operatorname*{arg\,min}_{y}\max_{y^{\prime}\in z}\|y-y^{\prime}\|^{2}$ is the Chebyshev center of the set $z$. Denote the Chebyshev radius of set $z$, by $r_{c}(z):=\min_{y}\max_{y^{\prime}\in z}\|y-y^{\prime}\|^{2}$. Let $\xi_{q_{A}}=g_{q_{A}}(s_{q_{A}}(\xi))$. The following lower bound holds for the average distortion of the abstraction:
@@
Prop. 5.1 suggests that the average distortion of an abstraction is lower bounded by the expected distortion of a particular encoder-decoder pair (the one outputting the Chebyshev centers of the abstractions outputs) plus a term depending on the size of the abstraction’s outputs. Employing Prop. 5.1, in Theorem 5.2 below, we derive fundamental lower bounds on $D_{abs}(R)$ and $R_{abs}(D)$, by lower-bounding each of the two terms in the right-hand side of (8) separately, over all abstractions with the same rate (or the same expected distortion). The first term in the right-hand side of (8) can be lower bounded as in Thm. 2.1, being the expected distortion of an encoder-decoder pair with the same rate as the abstraction. To bound the second term, we observe that the abstraction’s outputs $\Omega_{A}$ define an nlnl-dimensional cover [^6] This cover is precisely $\mathcal{Z}:=\{Z:Z=g_{A}(s_{A}(x_{0})),x_{0}\in\mathcal{X}\}$ and note that $s_{A}(x)$ takes values in the set $|\mathcal{Y}|$. Thus $|\mathcal{Z}|=|\mathcal{Y}|$. of ${\mathcal{B}}_{l}^{S}$, and the cover’s size is equal to the abstraction’s size; the bound is then obtained by lower-bounding over all possible nlnl-dimensional covers of ${\mathcal{B}}_{l}^{S}$, using geometric measure theory (see Lemma 8.1). For an illustrative example of the above, see Fig. 3.
Figure 3: Consider the dynamical system $x^{+}=x^{2}$ with state-space $\mathcal{X}=[0,1]$. The parabola depicts the set of trajectories ${\mathcal{B}}_{2}^{S}$, embedded in $0,1]^{2}$. Consider an abstraction AA with associated partition sets $Y_{i}=[0.2(i-1),\,0.2i)$ for $i=1,\dots,4$ and $Y_{5}=[0.8,1]$. The abstraction transitions, thus, are $\underset{A}{\rightarrow}=\{(Y_{1},Y_{1}),(Y_{2},Y_{1}),(Y_{3},Y_{1}),(Y_{3},Y_{2}),(Y_{4},Y_{2}),(Y_{4},Y_{3}),\allowbreak(Y_{4},Y_{4}),(Y_{5},Y_{4}),(Y_{5},Y_{5})\}$. The colored rectangles represent the abstraction outputs $\Omega_{A}$, depending on the initial condition $\xi_{0}$. For example, if $\xi_{0}\in Y_{4}$, then the abstraction output $\Omega_{A}$ is the green rectangle $\{(x_{0},x_{1}):\,x_{0}\in Y_{4},x_{1}\in Y_{3}\cup Y_{4}\cup Y_{5}\}$. Observe how the abstraction’s outputs define a 22-dimensional cover of the curve ${\mathcal{B}}_{2}^{S}$. The dots represent the Chebyshev centers for each different abstraction output, and the circles are the corresponding Chebyshev balls. E.g, when $\xi_{0}\in Y_{4}$, we have $x_{c}(\Omega_{A})=(.5\,\,.5)$ and $r_{c}(\Omega_{A})=0.3\sqrt{2}$. The abstraction’s expected distortion is lower bounded as per (8).
@@
Consider a dynamical system $x^{+}=f(x)$ with transition system representation $S=(\mathcal{X},\underset{S}{\rightarrow})$, and let Assumption 1 hold. Let $\xi\in{\mathcal{B}}_{l}^{S}$ be a trajectory of $S$, with $\xi_{0}\sim p_{\xi_{0}}:\mathcal{X}\to\mathbb{R}_{+}$. Assume that:
@@
The average distortion of any abstraction AA with partition size $|\mathcal{Y}|\leq e^{R}$, where $R>0$, is lower bounded as follows
@@
Notice that a valid lower bound is obtained for any value of $s\in(1,\infty]$ in the right-hand side of (9); maximization over $s$ provides the tightest bound. In the numerical examples in Section 6, we compute the bound for multiple values of $s$. Further, one may recover a lower bound on $R_{abs}(D)$ numerically, by fixing $D$ in the left-hand side of (9) and solving numerically for $R$ (as the right-hand side is a decreasing function of $R$, this is trivially computed by, e.g., bisection methods). The above theorem, thus, provides fundamental limits on the accuracy-size tradeoff, or the scalability, of abstractions, for given dynamics $x^{+}=f(x)$.
@@
The first assumption of Thm. 5.2, requiring ${\mathcal{B}}^{S}_{l}$ to be a union of smooth manifolds, is satisfied whenever the dynamics $f$ is piecewise continuously-differentiable. The second assumption is satisfied whenever $f$ is piecewise continuous and the initial condition distribution $p_{\xi_{0}}$ is such that $\mu_{\xi_{0}}\ll{\mathcal{L}}^{n}$, where ${\mathcal{L}}^{n}$ the Lebesgue measure.
In the coming section, we provide: a) closed-form expressions on $h(\xi)$, $h_{s}(\xi)$ and $c_{{\mathcal{B}}_{l}^{S}}$, for certain classes of dynamics $x^{+}=f(x)$, and b) an interpretation on how the complexity of the dynamics and the time-horizon $l$ affect the fundamental lower bound (9) in Thm. 5.2.
@@
where $J_{b_{l}}$ denotes the Jacobian matrix of $b_{l}$.
@@
As the partition size $|\mathcal{Y}|$ grows large, the Chebyshev balls of the abstraction outputs (c.f. Prop. 5.1 and Lemma 8.1) become small. Hence, in the case of smooth $f$, their intersection with the manifold approaches the case of an affine system, with $c_{{\mathcal{B}}_{l}^{S}}\leq v_{n}.$ Similarly, in the piecewise affine case, for sufficiently small balls – at least $|\mathcal{Y}|\geq M^{l}$ –, these can be chosen to intersect with at most one piece each. Thus, to reduce conservatism of the bound in such high-rate cases, one can inspect the lower bound of Thm. 2.1 by using $c_{{\mathcal{B}}_{l}^{S}}=v_{n}.$ We demonstrate this in the numerical examples in Section 6.
@@
Consider a system $x^{+}=f(x)$ whose origin is exponentially stable on a given compact set in $\mathbb{R}^{n}$. Then, there is a Lyapunov function $V:\mathbb{R}^{n}\mapsto\mathbb{R}_{+}$ satisfying $V(x)\geq\frac{1}{r}\|x\|^{2}$ for a given $r>0$ and, for all $x$ s.t. $V(x)\leq 1$ (w.l.o.g.), $V(f(x))\leq aV(x),$ with $a\in[0,1).$ This allows us to create an abstraction AA with the associated partition $Y_{i}=\{x\in\mathbb{R}^{n}\mid a^{i}<V(x)\leq a^{i-1}\}$ for i=1,,N,i=1,...,N, and $Y_{N+1}=\{x\in\mathbb{R}^{n}\mid V(x)\leq a^{N}\}$; and transitions $Y_{i}\xrightarrow[A]{}Y_{j}$ if and only if $j<i$ or $j=i=N+1$. The abstraction encapsulates the fact that, after NN steps or less, all trajectories reach the sublevel set $V(x)\leq a^{N}$. We get that $2r$ and $2a^{n}r$ are overapproximations of the diameters of $Y_{i},~~i\leq N,$ and $Y_{N+1},$ respectively. Then, recalling the distortion (7), for any trajectory $\xi$ of the system, $l>N,$
@@
which can be made arbitrarily small by suitable choice of NN.
@@
For a Schur LTI system, we have $c_{{\mathcal{B}}^{S}_{l}}\leq v_{n}$ and
@@
For the simple system $x^{+}=x$, we have that $\det(J_{b_{l}}(x)^{\mkern-1.5mu\mathrm{T}}\!J_{b_{l}}(x))=\det(lI)=l^{n}$ and, by Prop. 5.3, $h(\xi)=h(x_{0})+\frac{n\log l}{2}$ and
@@ 6.1 The doubling map
Consider the doubling map from Example 2.1. For any trajectory length $l$, its behavior ${\mathcal{B}}_{l}^{S}$ is composed by $2^{l-1}$ line segments in $\mathbb{R}^{l}$ described by $(x_{0},2x_{0},4x_{0},...,2^{l-1}x_{0})\bmod 1,$ uniformly distributed with $p_{\xi}(\xi)=1/\sqrt{1+4+...+4^{l-1}}=\sqrt{3/(4^{l}-1)},$ giving $h(\xi)=h_{s}(\xi)=\frac{1}{2}(\log(4^{l}-1)-\log{3})$ for all $s\in(1,\infty].$ Using Prop. 5.4 for piecewise affine systems, we obtain $c_{{\mathcal{B}}_{l}^{S}}\leq\sqrt{(4^{l}-1)/3}v_{n},$ enabling us to compute the lower bound in Theorem 5.2. In light of Remark 5, we also determine the high-rate lower bound by picking $c_{{\mathcal{B}}_{l}^{S}}=v_{1}=2.$ The lower-bound curves can be seen in Fig. 4 for $s=2$ and $s=\infty$. It is apparent that the tightest bound is obtained with $s=\infty$ and $c_{{\mathcal{B}}_{l}^{S}}=v_{1}.$ In this case, the bound is consistently half of that of the optimal distortion $D_{abs}(R)$, which is remarkably close. As a comparison, the standard Shannon lower bound is  1.42 times smaller than the optimal quantizer distortion of a uniform random variable in $\mathbb{R}^{1}$, in the standard source-coding setting.
@@
Let us explain how we were able to compute the actual optimal achievable abstraction distortion. Following the reasoning in Section 5, we first build an optimal cover for ${\mathcal{B}}_{l}^{S}$ (afterwards, we show that this optimal cover admits a distortion that is equal to that of a specific abstraction with the same rate, and thus its rate-distortion curve is optimal, among all abstractions). For a given $l$, consider $R=\log(k2^{l-1})$, where $k$ is an arbitrary natural number. Since all segments are equiprobable and congruent, and probability is uniform among them, the optimal partition of ${\mathcal{B}}_{l}^{S}$ is obtained by cutting each of the $2^{l-1}$ segments in $k$ equal pieces. The expected error between a trajectory $\xi$ and the Chebyshev center of its corresponding piece is $\mathbb{E}[\|\xi-\xi_{q_{A}}\|^{2}]=\frac{1-4^{-l}}{9k^{2}},$ which is obtained by computing the squared length of each segment, $L^{2}=1/(2^{l-1}k)^{2}(1+2^{2}+...+(2^{l-1})^{2})=(4^{l}-1)/(3k^{2}4^{l-1})=4(1-4^{-l})/3k^{2},$ followed by using the variance of the uniform distribution, giving $L^{2}/12$. To determine the corresponding abstraction lower bound, we use the distortion $d$ in (7) on the aforementioned pieces. For one dimensional line segments, $\max_{\xi^{\prime}\in\Omega_{A}(\xi)}\|\xi-\xi^{\prime}\|^{2}=(\|\xi-\xi_{q_{A}}\|+L/2)^{2}$. Its expected value is thus the second moment of a uniform from $L/2$ to $L$. Using $\mathbb{E}[X^{2}]=\mathrm{Var}(X)+\mathbb{E}[X]^{2}=L^{2}/48+9L^{2}/16=7L^{2}/12,$ we obtain $D_{cover}(R)=\frac{7}{l}4^{l-2}(4^{l}-1)e^{-2R},$ where we used $R=\log(k2^{l-1})$, and $D_{cover}(R)$ is the optimal distortion among all $l$-dimensional covers of ${\mathcal{B}}_{l}^{S}$.
@@
Notice that, in general, $D_{abs}(R)\geq D_{cover}(R)$, as abstractions are covers. However, the optimal cover built above determines an abstraction AA that gives the same distortion, and thus we have $D_{abs}(R)=D_{cover}(R)$. First, $\mathcal{Y}$ is the uniform grid with segments of length $1/k2^{l-1}$. Each trajectory of the abstraction is a sequence of segments of lengths $1/k2^{l-1},1/k2^{l-2},...,1/k$, thus giving a box in $\mathbb{R}^{l}$ containing any related trajectory $\xi$. For each box, the set of related trajectories is precisely a diagonal of the box. As such, the furthest edge along the diagonal is again a solution to $\arg\sup_{\xi^{\prime}\in\Omega_{A}(\xi)}\|\xi-\xi^{\prime}\|^{2}$. hence, the abstraction has the same distortion as the optimal cover. The above reasoning is illustrated in Fig. 5

### v3

@@ 5.1 A fundamental limit on abstracting dynamical systems
Consider a dynamical system $x^{+}=f(x)$ with transition system representation $S=(\mathcal{X},\underset{S}{\rightarrow})$, and let Assumption 1 hold. Let $\xi\in{\mathcal{B}}_{l}^{S}$ be a trajectory of $S$, with $\xi_{0}\sim p_{\xi_{0}}:\mathcal{X}\to\mathbb{R}_{+}$. Consider a measurable partition $\mathcal{Y}$ of $\mathcal{X}$ and an associated abstraction $A$, and let $\Omega_{A}=g_{A}(s_{A}(\xi))$, where $s,g$ are given by (4) and (5). Consider an encoder-decoder pair $(s_{q_{A}},g_{q_{A}})$, where $s_{q_{A}}(\xi)=g_{A}(s_{A}(\xi))$ and $g_{q_{A}}(z)=x_{c}(z)$, where $x_{c}(z):=\operatorname*{arg\,min}_{y}\max_{y^{\prime}\in z}\|y-y^{\prime}\|^{2}$ is the Chebyshev center of the set $z$. Denote the Chebyshev radius of set $z$, by $r_{c}(z):=\min_{y}\max_{y^{\prime}\in z}\|y-y^{\prime}\|^{2}$. Let $\xi_{q_{A}}=g_{q_{A}}(s_{q_{A}}(\xi))$. The following lower bound holds for the average distortion of the abstraction:
@@
Prop. 5.1 suggests that the average distortion of an abstraction is lower bounded by the expected distortion of a particular encoder-decoder pair (the one outputting the Chebyshev centers of the abstractions outputs) plus a term depending on the size of the abstraction’s outputs. Employing Prop. 5.1, in Theorem 5.2 below, we derive a fundamental lower bound on $D_{abs}(R)$, by lower-bounding each of the two terms in the right-hand side of (7) separately, over all abstractions with the same rate. The first term in the right-hand side of (7) can be lower bounded as in Thm. 2.1, being the expected distortion of an encoder-decoder pair with the same rate as the abstraction. To bound the second term, we observe that the abstraction’s outputs $\Omega_{A}$ define an $nl$-dimensional cover [^6] This cover is precisely $\mathcal{Z}:=\{Z:Z=g_{A}(s_{A}(x_{0})),x_{0}\in\mathcal{X}\}$ and note that $s_{A}(x)$ takes values in the set $|\mathcal{Y}|$. Thus $|\mathcal{Z}|=|\mathcal{Y}|$. of ${\mathcal{B}}_{l}^{S}$, and the cover’s size is equal to the abstraction’s size; the bound is then obtained by lower-bounding over all possible $nl$-dimensional covers of ${\mathcal{B}}_{l}^{S}$, using geometric measure theory (see Lemma 8.1). For an illustrative example of the above, see Fig. 3.
Figure 3: Consider the dynamical system $x^{+}=x^{2}$ with state-space $\mathcal{X}=[0,1]$. The parabola depicts the set of trajectories ${\mathcal{B}}_{2}^{S}$, embedded in $0,1]^{2}$. Consider an abstraction $A$ with associated partition sets $Y_{i}=[0.2(i-1),\,0.2i)$ for $i=1,\dots,4$ and $Y_{5}=[0.8,1]$. The abstraction transitions, thus, are $\underset{A}{\rightarrow}=\{(Y_{1},Y_{1}),(Y_{2},Y_{1}),(Y_{3},Y_{1}),(Y_{3},Y_{2}),(Y_{4},Y_{2}),(Y_{4},Y_{3}),\allowbreak(Y_{4},Y_{4}),(Y_{5},Y_{4}),(Y_{5},Y_{5})\}$. The colored rectangles represent the abstraction outputs $\Omega_{A}$, depending on the initial condition $\xi_{0}$. For example, if $\xi_{0}\in Y_{4}$, then the abstraction output $\Omega_{A}$ is the green rectangle $\{(x_{0},x_{1}):\,x_{0}\in Y_{4},x_{1}\in Y_{3}\cup Y_{4}\cup Y_{5}\}$. Observe how the abstraction’s outputs define a 22-dimensional cover of the curve ${\mathcal{B}}_{2}^{S}$. The dots represent the Chebyshev centers for each different abstraction output, and the circles are the corresponding Chebyshev balls. E.g, when $\xi_{0}\in Y_{4}$, we have $x_{c}(\Omega_{A})=(.5\,\,.5)$ and $r_{c}(\Omega_{A})=0.3\sqrt{2}$. The abstraction’s expected distortion is lower bounded as per (7).
@@
Consider a dynamical system $x^{+}=f(x)$ with transition-system representation $S=(\mathcal{X},\underset{S}{\rightarrow})$, and let Assumption 1 hold. Let $\xi\in{\mathcal{B}}_{l}^{S}$ be a trajectory of $S$, with $\xi_{0}\sim p_{\xi_{0}}:\mathcal{X}\to\mathbb{R}_{+}$. Assume that:
@@
The average distortion of any abstraction $A$ with partition size $|\mathcal{Y}|\leq e^{R}$, where $R>0$, is lower bounded as follows
@@
Notice that a valid lower bound on $D_{abs}(R)$ is obtained for any value of $s\in(1,\infty]$ in the right-hand side of (8); maximization over $s$ provides the tightest bound. In the numerical examples in Section 6, we compute the bound for multiple values of $s$. Further, one may recover a lower bound on $R_{abs}(D)$ numerically, by fixing $D$ in the left-hand side of (8) and solving numerically for $R$ (as the right-hand side is a decreasing function of $R$, this is trivially computed by, e.g., bisection methods). Nonetheless, by specifically picking $s=\infty$, one may analytically invert the bound (8) to obtain a lower bound on $R_{abs}(D)$, as the following corollary shows.
Corollary 5.3 (to Thm. 5.2).
Under the assumptions of Thm. 5.2, the minimum abstraction size, for an upper threshold $D$ on abstraction distortion, is lower bounded as:
$R_{abs}(D)\geq$ $\frac{n}{2}\log\!\Big(\frac{n}{2lc_{{\mathcal{B}}_{l}^{S}}^{2/n}}\frac{e^{\frac{2}{n}h(\xi)-1}}{\Gamma(1+n/2)^{2/n}}+e^{\frac{2}{n}h_{\infty}(\xi)}\Big)$ (9)
$-\frac{n}{2}\log(D).$
Proof Sketch.
Pick $s=\infty$ in (8), fix the right-hand side equal to $D$, and solve for $R$. ∎
Hence, Thm 5.2 and Cor. 5.3 provide fundamental limits on the accuracy-size tradeoff, or the scalability, of abstractions, for given dynamics $x^{+}=f(x)$.
@@
The first assumption of Thm. 5.2, requiring ${\mathcal{B}}^{S}_{l}$ to be a union of smooth manifolds, is satisfied whenever the dynamics $f$ is piecewise continuously-differentiable. The second assumption is satisfied whenever $f$ is piecewise continuous and the initial condition distribution $p_{\xi_{0}}$ is such that $\mu_{\xi_{0}}\ll{\mathcal{L}}^{n}$, where ${\mathcal{L}}^{n}$ the Lebesgue measure in $\mathbb{R}^{n}$.
In the coming section, we provide: a) closed-form expressions on $h(\xi)$, $h_{s}(\xi)$ and $c_{{\mathcal{B}}_{l}^{S}}$, for certain classes of dynamics $x^{+}=f(x)$, and b) an interpretation on how the complexity of the dynamics and the time-horizon $l$ affect the fundamental lower bound (8) in Thm. 5.2. Afterwards, in Section 5.3, we show how the curse of dimensionality arises, through a relaxed version of the bounds of Thm. 5.2 and Cor. 5.3.
@@
where $J_{b_{l}}$ denotes the Jacobian matrix of $b_{l}$. Moreover, $h(\xi)\geq h(\xi_{0})$ and $h_{s}(\xi)\geq h_{s}(\xi_{0})$ for any $s\in[1,\infty).$
@@
As the partition size $|\mathcal{Y}|$ grows large, the Chebyshev balls of the abstraction outputs (c.f. Prop. 5.1 and Lemma 8.1) become small. Hence, in the case of smooth $f$, their intersection with the manifold approaches the case of an affine system, with $c_{{\mathcal{B}}_{l}^{S}}\leq v_{n}.$ Similarly, in the piecewise affine case, for sufficiently small balls – at least $|\mathcal{Y}|\geq M^{l}$ –, these can be chosen to intersect with at most one piece each. Thus, to reduce conservatism of the bound in such high-rate cases, one can inspect the lower bound of Thm. 5.2 by using $c_{{\mathcal{B}}_{l}^{S}}=v_{n}.$ We demonstrate this in the numerical examples in Section 6.
@@
Consider a system $x^{+}=f(x)$ whose origin is exponentially stable on a given compact set in $\mathbb{R}^{n}$. Then, there is a Lyapunov function $V:\mathbb{R}^{n}\mapsto\mathbb{R}_{+}$ satisfying $V(x)\geq\frac{1}{r}\|x\|^{2}$ for a given $r>0$ and, for all $x$ s.t. $V(x)\leq 1$ (w.l.o.g.), $V(f(x))\leq aV(x),$ with $a\in[0,1).$ This allows us to create an abstraction $A$ with the associated partition $Y_{i}=\{x\in\mathbb{R}^{n}\mid a^{i}<V(x)\leq a^{i-1}\}$ for i=1,,N,i=1,...,N, and $Y_{N+1}=\{x\in\mathbb{R}^{n}\mid V(x)\leq a^{N}\}$; and transitions $Y_{i}\xrightarrow[A]{}Y_{j}$ if and only if $j<i$ or $j=i=N+1$. The abstraction encapsulates the fact that, after $N$ steps or less, all trajectories reach the sublevel set $V(x)\leq a^{N}$. We get that $2r$ and $2a^{n}r$ are overapproximations of the diameters of $Y_{i},~~i\leq N,$ and $Y_{N+1},$ respectively. Then, recalling the distortion (6), for any trajectory $\xi$ of the system, $l>N,$
@@
which can be made arbitrarily small by suitable choice of $N$.
@@
For an LTI system $x^{+}=Ax$, where the matrix $A$ is Schur-stable (its spectral norm is less than 1), we have $c_{{\mathcal{B}}^{S}_{l}}\leq v_{n}$ and
@@
For the simple system $x^{+}=x$, we have that $\det(J_{b_{l}}(x)^{\mkern-1.5mu\mathrm{T}}\!J_{b_{l}}(x))=\det(lI)=l^{n}$ and, by Prop. 5.4, $h(\xi)=h(x_{0})+\frac{n\log l}{2}$ and
@@
5.3 A relaxed lower bound and the curse of dimensionality
We now derive relaxed lower bounds that do not depend on the dynamics, and thus do not require computing trajectory entropies, as the following corollary to Thm. 5.2 shows.
Corollary 5.6 (Relaxed lower bound for abstractions).
Consider a dynamical system $x^{+}=f(x)$ under the assumptions of Thm. 5.2, with $\xi_{0}\sim p_{\xi_{0}}:\mathcal{X}\to\mathbb{R}_{+}$. Assume that $f$ is either (piecewise) affine or Lipschitz continuous with constant $L$. The average distortion of any abstraction $A$ with partition size $|\mathcal{Y}|\leq e^{R}$, where $R>0$, is lower bounded as follows:
$D_{abs}(R)\geq\,K(l)\bigg($ $\frac{n}{2}\Big(\frac{e^{-R+h(\xi_{0})-n/2}}{\Gamma(1+n/2)}\Big)^{2/n}$ (14)
$+\max_{s\in(1,\infty]}e^{\frac{2}{n}(-\frac{s}{s-1}R+h_{s}(\xi_{0}))}\bigg),$
where
  1. 1.
    $K(l)=\frac{1-L^{2}}{l{v_{n}}^{2/n}}$, if $f$ is Lipschitz with $L<1$;
  2. 2.
    $K(l)=\frac{1-L^{2}}{l^{2}{v_{n}}^{2/n}},$ if $f$ is Lipschitz with $L=1$;
  3. 3.
    $K(l)=\frac{1-L^{2}}{lL^{2l}{v_{n}}^{2/n}}$ if $f$ is Lipschitz with $L>1$;
  4. 4.
    $K(l)=\frac{1}{lM^{2l/n}{v_{n}}^{2/n}},$ if $f$ is piecewise affine with $M$ pieces.
Moreover, picking $s=\infty$ gives the following lower bound on $R_{abs}(D):$
$R_{abs}(D)\geq$ $\frac{n}{2}\log\!\Big(K(l)\frac{n}{2}\frac{e^{\frac{2}{n}h(\xi_{0})-1}}{\Gamma(1+n/2)^{2/n}}+e^{\frac{2}{n}h_{\infty}(\xi_{0})}\Big)$ (15)
$-\frac{n}{2}\log(D).$
Poof Sketch.
The result follows from Thm. 5.2, by combining Propositions 5.4 and 5.5 and using the fact that $\sum_{i=0}^{k-1}a^{i}=k$ for $a=1$, and otherwise $\sum_{i=0}^{k-1}a^{i}=(a^{k}-1)/(a-1)\leq\max(1,a^{k})/|a-1|$. ∎
While more conservative than Thm. 5.2, Cor. 5.6 confirms the curse of dimensionality for abstractions: for any system and fixed finite $l$, (15) shows that the size of the partition $e^{R}$ fundamentally increases exponentially with $n$ to achieve a prescribed distortion.
@@ 6.1 The doubling map
Consider the doubling map from Example 2.1. For any trajectory length $l$, its behavior ${\mathcal{B}}_{l}^{S}$ is composed by $2^{l-1}$ line segments in $\mathbb{R}^{l}$ described by $(x_{0},2x_{0},4x_{0},...,2^{l-1}x_{0})\bmod 1,$ uniformly distributed with $p_{\xi}(\xi)=1/\sqrt{1+4+...+4^{l-1}}=\sqrt{3/(4^{l}-1)},$ giving $h(\xi)=h_{s}(\xi)=\frac{1}{2}(\log(4^{l}-1)-\log{3})$ for all $s\in(1,\infty].$ Using Prop. 5.5 for piecewise affine systems, we obtain $c_{{\mathcal{B}}_{l}^{S}}\leq\sqrt{(4^{l}-1)/3}v_{n},$ enabling us to compute the lower bound in Theorem 5.2. In light of Remark 6, we also determine the high-rate lower bound by picking $c_{{\mathcal{B}}_{l}^{S}}=v_{1}=2.$ The lower-bound curves can be seen in Fig. 4 for $s=2$ and $s=\infty$. It is apparent that the tightest bound is obtained with $s=\infty$ and $c_{{\mathcal{B}}_{l}^{S}}=v_{1}.$ In this case, the bound is consistently half of that of the optimal distortion $D_{abs}(R)$, which is remarkably close. As a comparison, the standard Shannon lower bound is 1.42 times smaller than the optimal quantizer distortion of a uniform random variable in $\mathbb{R}^{1}$, in the standard source-coding setting with Euclidean distortion.
@@
Let us explain how we were able to compute the actual optimal achievable abstraction distortion. Following the reasoning in Section 5, we first build an optimal cover for ${\mathcal{B}}_{l}^{S}$ (afterwards, we show that this optimal cover admits a distortion that is equal to that of a specific abstraction with the same rate, and thus its rate-distortion curve is optimal, among all abstractions). For a given $l$, consider $R=\log(k2^{l-1})$, where $k$ is an arbitrary natural number. Since all segments are equiprobable and congruent, and probability is uniform among them, the optimal cover of ${\mathcal{B}}_{l}^{S}$ is obtained by cutting each of the $2^{l-1}$ segments in $k$ equal pieces. [^8] For uniform distributions, with convex and symmetric distortion functions as the one in (6), optimal covers (or quantizers) are uniform [32]. The cover is then the collection $M=\{C_{i}\}_{i=1}^{2^{l-1}k}$ of all these pieces—see Fig. 5 for an example with $l=3$ and $k=2$. Now let $C_{\xi}\in M$ denote the piece including the trajectory $\xi$. The expected error between $\xi$ and the Chebyshev center of $C_{\xi}$ is $\mathbb{E}[\|\xi-x_{c}(C_{\xi})\|^{2}]=\frac{1-4^{-l}}{9k^{2}},$ which is obtained by computing the squared length of each segment, $L^{2}=1/(2^{l-1}k)^{2}(1+2^{2}+...+(2^{l-1})^{2})=(4^{l}-1)/(3k^{2}4^{l-1})=4(1-4^{-l})/3k^{2},$ followed by using the variance of the uniform distribution, giving $L^{2}/12$. Then, for the cover’s set-based distortion, we have $d(\xi,C_{\xi})=\frac{1}{l}\max_{\xi^{\prime}\in C_{\xi}}\|\xi-\xi^{\prime}\|^{2}=\frac{1}{l}(\|\xi-x_{c}(C_{\xi})\|+L/2)^{2}$, since $C_{\xi}$ is a straight-line segment. We, thus have $\mathbb{E}[d(\xi,C_{\xi})]=\frac{1}{l}\mathbb{E}[(\|\xi-x_{c}(C_{\xi})\|+L/2)^{2}]=7L^{2}/(12l)$, where we used that $X\coloneqq\|\xi-x_{c}(C_{\xi})\|+L/2$ is a uniform random variable in $[L/2,L]$, and thus $E[X^{2}]=\mathrm{Var}(X)+\mathbb{E}[X]^{2}=L^{2}/48+9L^{2}/16=7L^{2}/12$. As $M$ is the optimal cover, we have $D_{cover}(R)=\mathbb{E}[d(\xi,C_{\xi})]=\frac{7}{l}4^{l-2}(4^{l}-1)e^{-2R},$ where we used $R=\log(k2^{l-1})$, and $D_{cover}(R)$ is the optimal distortion among all $nl$-dimensional covers of ${\mathcal{B}}_{l}^{S}$.
@@
Notice that, in general, $D_{abs}(R)\geq D_{cover}(R)$, as behaviors of abstractions are covers. However, the optimal cover built above admits an abstraction $A$ that gives the same distortion, and thus we have $D_{abs}(R)=D_{cover}(R)$. First, let $\mathcal{Y}$ be the uniform grid of $[0,1]$ with segments of length $1/k2^{l-1}$. Each trajectory of the abstraction is a sequence of segments of lengths $1/k2^{l-1},1/k2^{l-2},...,1/k$; thus, for any given trajectory $\xi$, the abstraction output $\Omega_{A}(\xi)$ is a box in $\mathbb{R}^{l}$ containing any related trajectory $\xi$ (see Fig. 5). In fact, $\xi$ lies on one of the diagonals of the box $\Omega_{A}(\xi)$, and this diagonal is precisely $C_{\xi}$. We thus have $\arg\sup_{\xi^{\prime}\in\Omega_{A}(\xi)}\|\xi-\xi^{\prime}\|^{2}=\arg\sup_{\xi^{\prime}\in C_{\xi}}\|\xi-\xi^{\prime}\|^{2}$, and hence $d(\xi,\Omega_{A}(\xi))=d(\xi,C_{\xi})$. Consequently, the abstraction has the same distortion as the optimal cover.